A first-order reaction has a rate constant of 7.5 × 10-3/s. The time required for the reaction to be 60% complete is

3.8 × 10-3 s

6.9 × 10-3 s

68 s

120 s

130 s
Answer #1

for a 1st order reaction -dA/dt= K[A], where K is rate constant and [A]= concentration of reactant

when integrated noting that at t=0 [A]=[A]o and at t=t, [A] =[A

lnA= lnAO-Kt (1)

but conversion XA= 1- [A]/[A]o

hence equation 1 in terms of XA becomes -ln(1-XA)= Kt

given XA=0.6, -ln(1-0.6)= 7.5*10-3*t, t= 122 seconds ( D is correct)